# CLF-RL: Control Lyapunov Function Guided Reinforcement Learning

### v1

@@ III-A Reference Trajectory Generation
$\text{Physical constraints (e.g., torque, joint limits)},$
$y(q(t))=B(\tau(t);\alpha).$
Here, $y(q)$ denotes the controlled outputs, and $B(\tau;\alpha)$ is the Bézier-parameterized desired trajectory. The phase variable $\tau(t)=\phi(q,t)$ indicate the progression along the trajectory. In our implementation, $\tau$ is time-based, however, a common alternative is a state-based phase variable, such as the horizontal hip displacement, which ensures monotonicity. This setup guarantees that the Bézier curve is evaluated at the correct phase value during optimization.
The optimized trajectory specifies desired joint or end-effector positions and velocities as functions of phase over a single step with one stance leg. Once a single step is solved, the solution is symmetrically remapped to produce a full gait cycle with alternating stance legs, yielding reference trajectories for the complete walking motion.
III-B CLF-Based Reward Terms
We denote the output error between the current state and reference trajectory as
$\eta(t)=\begin{pmatrix}y^{d}(t)-y(q)\\ \dot{y}^{d}(t)-\dot{y}(q,\dot{q})\end{pmatrix}$
Based on this tracking error, we construct a control Lyapunov function as:
$V(\eta)=\eta^{\top}P\eta,$
This formulation implicitly assumes a double integrator reference model for the output dynamics, allowing us to define a Lyapunov function without explicitly computing Lie derivatives $L_{f}V$ or $L_{g}V$ for the full-order system. The matrix $P\succ 0$ is the unique solution to the Continuous-Time Algebraic Riccati Equation (CARE) corresponding to this linearized model.
To avoid explicitly computing $\dot{\eta}$ and $\dot{V}$, we approximate the Lyapunov derivative using finite differences:
$\dot{V}_{t}\approx\frac{V_{t+1}-V_{t}}{\Delta t},\quad\text{with}\quad V_{t}=V\bigl{(}\eta(t)\bigr{)}.$
We define a CLF tracking reward:
$r_{v}$ $=w_{v}\exp\left(-\frac{V_{t}}{\sigma_{v}}\right)$ (CLF Tracking)
where $\sigma_{v}=\mu_{\max}(P)\,\eta_{\max}^{2}$, with $\mu_{\max}(P)$ denoting the maximum eigenvalue of the CLF matrix $P$, $\eta_{\max}$ representing an empirical bound on the tracking error.
We consider two versions of the CLF decay condition: a smooth formulation using the hyperbolic tangent, and a clipped version:
$r_{\dot{v}}^{\tanh}$ $=w_{\dot{v}}\tanh\left(-\left(\dot{V}_{t}+\lambda V_{t}\right)\right)$ (CLF Decay: Tanh)
$r_{\dot{v}}^{\text{clip}}$ $=w_{\dot{v}}\cdot\max\left(\min\left(\frac{\dot{V}_{t}+\lambda V_{t}}{\sigma_{\dot{v}}},\,1\right),\,0\right),$ (CLF Decay: Clipped)
[…]
III-C Other Reward Terms
In addition to the CLF terms, we incorporate several auxiliary rewards commonly used in RL for legged system:
Stance Foot Holonomic Reward: To enforce stance-foot holonomic constraints, we define a combined reward:
@@ III-C Other Reward Terms
where $p_{\mathrm{st}}$ and $p_{\mathrm{st}}^{0}$ are the current and the initial stance foot positions when it enter the domain, $v_{\mathrm{st}}$ is the stance foot velocity, and $w_{\mathrm{hpos}},w_{\mathrm{hvel}}$ are weighting terms.
Regularization Term: To discourage excessive control effort, abrupt actions, and joint limit violations, we include a regularization reward:
@@
$-w_{q_{\text{limit}}}\left\|\max\left(0,q_{\min}-q_{t}\right)+\max\left(0,q_{t}-q_{\max}\right)\right\|_{1},$
@@
III-D Total Reward
The final reward used for policy training is the weighted sum of all components:
$R=\;r_{v}+r_{\dot{v}}+r_{\mathrm{hol}}+r_{\mathrm{reg}},$
@@ III-D Total Reward
This reward structure encodes both task performance and physical feasibility, allowing the policy to balance tracking, stability, and regularization objectives. By combining model-based references (e.g., from reduced-order planning or offline trajectory optimization) with CLF-inspired reward shaping and constraint-aware regularization, our framework embeds control-theoretic structure into the reinforcement learning process. This guides training toward stable, robust, and physically plausible locomotion, while preserving the flexibility and adaptability of model-free RL.
IV CLF-RL Implementation
@@
We demonstrate our framework on the 29-DoF Unitree G1 humanoid robot, using 21 actuated degrees of freedom for motion planning. The hand and wrist joints are fixed during training and controlled with zero desired position and velocity during deployment.
Reference Generation: We use the same set of end-effector and joint positions for both HZD- and H-LIP-based trajectories. The full reference output $y_{\alpha}^{d}(t)\in\mathbb{R}^{n_{y}}$, with $n_{y}=21$ is structured as follows:
$y_{\alpha}^{d}=\begin{pmatrix}p^{d}_{\mathrm{com}}&\phi^{d}_{\mathrm{pelvis}}&p^{d}_{\mathrm{sw}}&\theta^{d}_{\mathrm{sw}}&q^{d}_{\mathrm{shoulder}}&q^{d}_{\mathrm{elbow}}\end{pmatrix}^{\top}$,
where the terms represent the desired CoM position, pelvis orientation, swing foot position and orientation, and arm joint angles. To account for turning motions during training, we heuristically adjust the reference trajectory based on the commanded angular velocity. Specifically, we modify the yaw orientation of all end-effector frames to align with the integrated yaw derived from the commanded angular velocity over time.
The HZD optimization is computed offline using IPOPT [26] and Casadi [27]. Hard constraints are used to enforce dynamics, periodicity, virtual constraints, and step length among others. The optimization problem is formulated as a multiple-shooting problem over a single swing phase.
@@
Policy Structure: The policy receives a combination of proprioceptive and task-relevant inputs, including angular velocity, projected gravity, commanded linear and angular velocities, relative joint positions and velocities, the previous action, and a phase-based time encoding using $\sin\left(2\pi t/t_{\mathrm{period}}\right)$ and $\cos\left(2\pi t/t_{\mathrm{period}}\right)$. The stepping period $t_{\mathrm{period}}$ is set to $0.8\,\mathrm{s}$, corresponding to a full gait cycle (i.e., two steps). The policy outputs joint position commands relative to a default symmetric standing pose at 50 Hz. This default pose is fixed across all policies we trained.
Both the actor and critic networks use a fully connected feedforward architecture with hidden layer dimensions of [512, 256, 128], using the ELU (Exponential Linear Unit) activation function at each layer. To facilitate learning, the critic (value network) receives additional privileged information $o^{\mathrm{priv}}_{t}$ not accessible to the actor, such as stance and swing foot linear and angular velocities, reference trajectory positions and velocities, and binary contact state indicators. These inputs help stabilize value estimation and improve training efficiency.
Training Procedure: We use IsaacLab, a GPU-accelerated simulation framework built on top of NVIDIA Isaac Sim, with its development initiated from the Orbit framework [28]. For policy training, we use the PPO implementation from Robotics System Lab RL library [29].
All training use the same set of terrain settings, robot models, and domain randomization parameters. To improve policy robustness and facilitate transfer to hardware, we apply domain randomization to physical properties such as link masses, static and dynamic friction coefficients, and the center of mass position. Additionally, external perturbations are introduced by applying randomized velocity impulses in the $x$ and $y$ directions to the base link at fixed time intervals.
Training is conducted over a range of commanded velocities, with linear velocity $v_{x}\in[-0.75,0.75]$ and yaw rate $\omega_{z}\in[-0.5,0.5]$. Given a commanded velocity, $v^{d}$, the reference trajectory is either retrieved from the precomputed gait library by selecting the closest matching forward velocity in the case of HZD, or generated online using the analytic form of the H-LIP gait library, which provides closed-form solutions across the velocity space.
The baseline policy is trained with a manually designed reward composed of several heuristic terms. These include tracking of linear $x$, $y$ and angular yaw velocity, while penalizing vertical velocity, angular velocity about the $x$ and $y$ axes, joint accelerations, joint velocities, joint torques, and abrupt changes in actions. The reward also discourages deviations from an upright base orientation and enforces joint limit compliance. Additional terms encourage behaviors such as staying upright (“stay alive”), avoiding foot slip during contact, maintaining a nominal hip posture and torso height, ensuring adequate foot clearance, and regulating contact scheduling. Furthermore, the policy is penalized for deviations of the arm and torso joints from predefined reference configurations.
TABLE I: Reward weight coefficients used during training for both HZD-CLF and LIP-CLF.
@@
V Results
Figure 3: Tracking performance comparison between a policy trained using only the reference tracking reward ($r_{v}$) and one trained with both reference tracking and CLF decrease condition rewards ($r_{v}$ and $r_{\dot{v}}$). Overall, the HZD-CLF policy achieves better tracking performance
@@ V Results
Figure 4: Tracking performance with torso mass randomly displaced within a box of size $\pm[0.05\text{ (x)},0.05\text{ (y)},0.01\text{ (z)}]$m around the nominal location. Fifty displacements are uniformly sampled, and the resulting mean and standard deviation of performance are plotted. The CLF-RL policies demonstrate lower variability, indicating improved consistency and robustness across different mass configurations.
We evaluate three policy variants: a baseline trained with hand-tuned rewards, a H-LIP-based CLF shaped policy (LIP-CLF), and an HZD-based CLF-shaped policy (HZD-CLF). All policies are validated in both sim-to-sim transfer using MuJoCo [30] and on hardware (Fig. 6). The experiment video can be found [31]. The policy runs onboard on an additional laptop mounted to the front of the torso during deployment, which also added additional mass (0.616 kg) to the robot.
CLF Decay Condition (Sim): To evaluate whether reference tracking alone is sufficient, or if enforcing the CLF decay condition provides additional benefit, we conduct an ablation study. Fig. 3 compares two policies: one trained with only the tracking reward ($r_{v}$), and another with both the tracking and CLF decay rewards ($r_{v}$ and $r_{\dot{v}}$). Both policies track the commanded velocity $v_{x}^{d}$ reasonably well across varying commands. However, the HZD-CLF policy shows improved steady-state tracking, maintaining average velocity closer to the desired value in forward and backward walking phases. While the overall transient behavior remains similar, the inclusion of the CLF decay condition reduces residual tracking error and promotes more consistent steady-state performance.
@@
Mass Location Randomization (Sim): To assess robustness to modeling mismatches, we introduce random perturbations to the torso mass location to simulate sim-to-real discrepancies in inertial properties. Specifically, the torso center of mass is uniformly displaced within a box of size $\pm[0.05~\text{m}~(x),0.05~\text{m}~(y),0.01~\text{m}~(z)]$. For each policy, we sample 50 randomized configurations and evaluate the velocity tracking performance across episodes, reporting the mean and standard deviation. As shown in Fig. 4, both CLF-RL policies exhibit significantly lower performance variance compared to the baseline. This indicates improved robustness and consistency under structural uncertainties.
Added Mass Perturbation (Sim): To further evaluate robustness, we introduce an 8 kg payload to the torso and compare tracking performance across the three policy variants under a velocity ramp command. As shown in Fig. 5, the CLF-based RL policies maintain more accurate velocity tracking, demonstrating improved resilience to payload-induced dynamics changes.
@@
Figure 5: Robustness testing in simulation with different policies: HZD-CLF, LIP-CLF and a baseline RL policy are compared with an additional 8kg mass added to the torso. A two second ramp up to the maximum trained velocity is commanded. The steady-state mean of the velocities is plotted in a dashed line. The CLF-shaped policies show superior robustness to the baseline.
@@
Indoor Experiments (Hardware): We deploy all policies on the Unitree G1 robot in a controlled indoor environment. Fig. 6 shows all three policies operating successfully on hardware. To quantify velocity tracking performance and robustness, we use a motion capture system to record global position and orientation data. This data is used solely for evaluation and is not fed into the controller. To remain consistent with simulation analysis, we track the pelvis frame and compute local-frame velocity by finite-differencing the position data (recorded at 240 Hz) and aligning it with the robot’s yaw orientation. Fig. 7 demonstrates the improved robustness of the HZD-CLF policy relative to the baseline. To test this robustness, we attach a backpack to the robot and load it with either 1.765 kg or 3.55 kg of additional mass. The HZD-CLF policy maintains consistent tracking performance across all conditions, whereas the baseline policy exhibits significant degradation. In the heavier case, the baseline policy drifts off the walking course and fails to complete the test. These hardware results corroborate the simulation findings in Fig. 4, which show that CLF-based policies exhibit lower performance variance under structural perturbations.
@@
Figure 7: Quantitative hardware testing shows the difference in robustness of the baseline and HZD-CLF policies. Additional mass is added to a backpack on the back and the velocity tracking is compared. We can see that the HZD-CLF policy has effectively no change in performance with the additional mass. For the heavier mass on the baseline, the policy drifted so much as to exit the walking course and collide with a table, as indicated in the plot.
Outdoor Experiments (Hardware): To evaluate the HZD-CLF policy under real-world conditions, we deploy the robot outdoors across a variety of environments. As shown in Fig. 8, the test route includes diverse flat-ground surfaces such as concrete and tiled walkways, as well as mild slopes like ADA-compliant ramps. The robot traverses all terrain types within a single continuous 0.25-mile walking trial without any failures. The distance was determined by experimental design considerations, rather than reflecting any limitation of the policy.
@@
VI Conclusion
We presented CLF-RL, a structured reward shaping framework that integrates reference trajectory tracking with control Lyapunov functions (CLFs). By embedding CLF-based objectives directly into the reinforcement learning reward, our approach replaces heuristic reward design with a theoretically grounded stability metric. We showed that incorporating the CLF decay condition improves tracking performance over baseline tracking-only rewards. CLF-based policies exhibit lower variance across a range of perturbations in both simulation and hardware, demonstrating increased reliability.
Our framework is modular and extensible. We demonstrated its compatibility with both reduced-order (H-LIP) and full-order (HZD) reference generators, though other trajectory planners could be incorporated. While our training only used steady-state reference motions, the learned policies generalize well to transient motions—highlighting the framework’s flexibility. Finally, we validated the real-world robustness of the approach through extended outdoor trials, showing consistent and stable walking performance across varied flat-ground terrains.

### v2

@@ III-B CLF-Based Reward Design
III-C Auxiliary Rewards and Final Reward Composition
In addition to the CLF terms, we incorporate several auxiliary rewards commonly used in RL for legged systems. To enforce stance-foot holonomic constraints, we define a combined reward:
@@ III-C Auxiliary Rewards and Final Reward Composition
where $p_{\mathrm{st}}$ and $p_{\mathrm{st}}^{0}$ are the current and the initial stance foot positions when it enters the domain, $v_{\mathrm{st}}$ is the stance foot velocity, and $w_{\mathrm{hpos}}$ and $w_{\mathrm{hvel}}$ are weights.
To discourage excessive control effort, abrupt actions, and joint limit violations, we include a regularization reward:
@@
- $w_{q_{\text{limit}}}\left\|\max\left(0,q_{\min}-q_{t}\right)+\max\left(0,q_{t}-q_{\max}\right)\right\|_{1},$
@@
The final reward used for policy training is the weighted sum of all components: $R=\;r_{v}+r_{\dot{v}}+r_{\mathrm{hol}}+r_{\mathrm{reg}}.$ This reward structure encodes both task performance and physical feasibility, allowing the policy to balance tracking, stability, and regularization objectives. By combining model-based references (e.g., from reduced-order planning or offline trajectory optimization) with CLF-inspired reward shaping and constraint-aware regularization, our framework embeds control-theoretic structure into the reinforcement learning process. This guides training toward stable, robust, and physically plausible locomotion, while preserving the flexibility and adaptability of model-free RL.
III-D Implementation Details
@@
We demonstrate our framework on the 29-DoF Unitree G1 humanoid robot, planning over 21 actuated DoF. The hand and wrist joints are fixed during training and controlled with zero desired position and velocity during deployment.
Reference Generation: We use the same set of end-effector and joint positions for both HZD and H-LIP-based trajectories. The full reference output $y_{\alpha}^{d}(t)\in\mathbb{R}^{n_{y}}$, with $n_{y}=21$ is structured as follows: $y_{\alpha}^{d}=\begin{pmatrix}p^{d}_{\mathrm{com}}&\phi^{d}_{\mathrm{pelvis}}&p^{d}_{\mathrm{sw}}&\theta^{d}_{\mathrm{sw}}&q^{d}_{\mathrm{shoulder}}&q^{d}_{\mathrm{elbow}}\end{pmatrix}^{\top}$ where the terms represent the desired CoM position, pelvis orientation, swing foot pose, and arm joint angles. One can verify that this set of outputs lead to an invertible $\frac{\partial\bar{y}}{\partial q}$ that satisfy Prop 1. To account for turning during training, we heuristically adjust the reference trajectory based on the commanded angular velocity. Specifically, we modify the yaw orientation of all end-effector frames to align with the integrated yaw derived from the commanded angular velocity over time. The HZD optimization is computed offline using IPOPT [wachter_implementation_2006] and Casadi [andersson_casadi_2019]. Hard constraints are used to enforce dynamics, periodicity, virtual constraints, and step length among others. The optimization problem is formulated as a multiple-shooting problem over a single swing phase.
@@
Policy Structure: The policy receives a combination of proprioceptive and task-relevant inputs, including angular velocity, projected gravity, commanded linear and angular velocities, relative joint positions and velocities, the previous action, and a phase-based time encoding using $\sin\left(2\pi t/t_{\mathrm{period}}\right)$ and $\cos\left(2\pi t/t_{\mathrm{period}}\right)$. The stepping period $t_{\mathrm{period}}$ is set to $0.8\,\mathrm{s}$, corresponding to a full gait cycle (i.e., two steps). The policy outputs joint position commands relative to a default symmetric standing pose at 50 Hz. This default pose is fixed across all policies we trained. Both the actor and critic networks use a fully connected feedforward architecture with hidden layer dimensions of [512, 256, 128] and Exponential Linear Unit activation function. The critic receives additional privileged information $o^{\mathrm{priv}}_{t}$, such as stance and swing foot linear and angular velocities, reference trajectory positions and velocities, and binary contact state indicators.
TABLE I: Reward weights used for HZD-CLF and LIP-CLF.
@@
Training Procedure: We use IsaacLab, a GPU-accelerated simulation framework built on top of NVIDIA Isaac Sim, with its development initiated from the Orbit framework [mittal2023orbit]. For policy training, we use the PPO implementation from Robotics System Lab RL library [rudin2022learning].
All training, including baseline and CLF-based results, use the same set of terrain settings, robot models, and domain randomization parameters. To improve policy robustness and facilitate transfer to hardware, we apply domain randomization to physical properties such as link masses, static and dynamic friction coefficients, and the center of mass position. Additionally, external perturbations are introduced by applying randomized velocity impulses in the $x$ and $y$ directions to the base link at fixed time intervals. Training is conducted over a range of commanded velocities, with linear velocity $v_{x}\in[-0.75,0.75]$ and yaw rate $\omega_{z}\in[-0.5,0.5]$. Given a commanded velocity, $v^{d}$, the reference trajectory is either retrieved from the precomputed gait library by selecting the closest matching forward velocity in the case of HZD, or generated online using the analytic form of the H-LIP gait library, which provides closed-form solutions across the velocity space. For both versions, we used the same reward listed under Table I.
The custom baseline policy is trained using a manually designed reward function composed of multiple heuristic terms. These include tracking linear velocities in the $x$ and $y$ directions as well as angular yaw velocity, while penalizing vertical velocity, angular velocities about the $x$ and $y$ axes, joint accelerations, joint velocities, joint torques, and abrupt changes in actions. The reward also discourages deviations from an upright base orientation and enforces compliance with joint limits. Additional terms promote behaviors such as maintaining an upright posture, preventing foot slip during contact, preserving nominal hip and torso positions, ensuring sufficient foot clearance, and regulating contact timing. Finally, the policy is penalized for deviations of the arm and torso joints from predefined reference configurations.
IV Results
@@
Figure 3: Tracking performance with torso mass randomly displaced within a box of size $\pm[0.05\text{ (x)},0.05\text{ (y)},0.01\text{ (z)}]$m around the nominal location. Fifty displacements are uniformly sampled, and the resulting mean and standard deviation of performance for each policy are plotted.
TABLE II: Tracking Velocity Performance Comparison.
Coord. Method Vel. Err. Mean $\downarrow$ Vel. Err. Std $\downarrow$
Center of Mass (cm/s)
$x$ CLF (ours) 7.3 (-24.4%) 4.6 (-20.1%)
Tracking only 9.6 5.8
$y$ CLF (ours) 5.9 (-4.6%) 3.7 (-7.0%)
Tracking only 6.2 4.0
$z$ CLF (ours) 7.2 (-7.3%) 4.4 (-4.7%)
Tracking only 7.7 4.7
Pelvis Orientation (rad/s)
Roll CLF (ours) 0.141 (-4.7%) 0.084 (-9.4%)
Tracking only 0.148 0.093
Pitch CLF (ours) 0.312 (-6.8%) 0.125 (-2.6%)
Tracking only 0.335 0.128
Yaw CLF (ours) 0.319 (-8.5%) 0.165 (-4.8%)
Tracking only 0.348 0.173
Swing Ankle (cm/s)
$x$ CLF (ours) 17.2 (-18.4%) 11.3 (-18.6%)
Tracking only 21.1 13.9
$y$ CLF (ours) 9.7 (-13.3%) 7.4 (-10.4%)
Tracking only 11.2 8.3
$z$ CLF (ours) 21.2 (+3.3%) 7.4 (-0.9%)
Tracking only 20.5 7.4
Swing Ankle Orientation (rad/s)
Roll CLF (ours) 0.346 (-7.6%) 0.272 (+13.2%)
Tracking only 0.374 0.240
Pitch CLF (ours) 0.641 (+8.9%) 0.445 (+4.6%)
Tracking only 0.589 0.425
Yaw CLF (ours) 0.228 (-5.2%) 0.176 (-7.5%)
Tracking only 0.241 0.190
We evaluate four policy variants: a baseline with rewards from the Unitree open-source implementation [^1] https://tinyurl.com/e5nrjdb9 (Unitree), a baseline trained with hand-tuned custom rewards (Custom), a H-LIP-based CLF shaped policy (LIP-CLF), and an HZD-based CLF-shaped policy (HZD-CLF) [^2] Open source implementation is available at https://github.com/Zolkin1/robot_rl. We also attempted to use potential shaping [jeon2023benchmarking] to replace the tracking reward, but the policy never learned to walk. We hypothesize that potential shaping works better when some key rewards are left in their nominal form, but in our proposed framework there is only one tracking reward. The CLF-based policies were trained with many random seeds without any noticeable difference in performance, demonstrating that the proposed RL training pipeline is robust. All policies are first validated in sim-to-sim transfer using MuJoCo [todorov2012mujoco]. Because the Unitree-baseline underperforms compared to the hand-tuned baseline (Fig. 3), only three policies are tested on hardware (Fig. 5). The policies run onboard on an additional laptop mounted to the front of the torso during deployment, which also added additional mass (0.616 kg) to the robot.
CLF Decay Condition: To assess whether reference tracking alone is sufficient or if enforcing the CLF decay condition offers additional benefits, we conduct an ablation study. Table II compares two policies: one trained with only the tracking reward ($r_{v}$) and another trained with both the tracking and CLF decay rewards ($r_{v}$ and $r_{\dot{v}}$). We simulate 200 instances of the robot under out-of-distribution domain randomization to evaluate robustness. The mean and standard deviation of the tracking errors are reported at steady state while walking at 0.75 m/s. The results show that including the CLF condition generally reduces both the mean tracking error and its variability, averaging about a 10% improvement in velocity tracking, and up to about 25%.
@@
Robustness evaluation: To assess robustness to modeling mismatches, we introduce random perturbations to the torso mass location to simulate sim-to-real discrepancies in inertial properties. Specifically, the torso center of mass is uniformly displaced within a box of size $\pm[0.05~\text{m}~(x),0.05~\text{m}~(y),0.01~\text{m}~(z)]$. For each policy, we sample 50 randomized configurations and evaluate the velocity tracking performance across episodes, reporting the mean and standard deviation. As shown in Fig. 3, both CLF-RL policies exhibit significantly lower performance variance compared to the hand-tuned baseline, with the unitree-baseline having the largest variance . This indicates improved robustness and consistency under structural uncertainties. To further evaluate robustness, we introduce an 8 kg payload to the torso and compare tracking performance across the three policy variants under a velocity ramp command. As shown in Fig. 4, the CLF-based RL policies maintain more accurate velocity tracking, demonstrating improved resilience to payload-induced dynamics changes.
@@
Figure 4: Robustness evaluation in simulation: HZD-CLF, LIP-CLF and the baseline policy are tested with an extra 8kg on the torso. A 2-second ramp commanded each controller up to the maximum training velocity. The steady-state mean velocities are indicated by dashed lines.
@@
Hardware Experiments: We deploy all the policies, except the Unitree baseline, on the Unitree G1 robot in a controlled indoor environment. Fig. 5 shows all three policies operating successfully on hardware. To quantify velocity tracking performance and robustness, we use a motion capture system to record global position and orientation data. This data is used solely for evaluation and is not fed into the controller. To remain consistent with simulation analysis, we track the pelvis frame and compute local-frame velocity by finite-differencing the position data (recorded at 240 Hz) and aligning it with the robot’s yaw orientation. Fig. 6 demonstrates the improved robustness of the HZD-CLF policy relative to the baseline. To test this robustness, we attach a backpack to the robot and load it with either 1.765 kg or 3.55 kg of additional mass. The HZD-CLF policy maintains consistent tracking performance across all conditions, whereas the baseline policy exhibits significant degradation. In the heavier case, the baseline policy drifts off the walking course and fails to complete the test. These hardware results corroborate the simulation findings in Fig. 3, which show that CLF-based policies exhibit lower performance variance under structural perturbations.
@@
Figure 6: Hardware testing with added backpack mass. HZD-CLF maintains good tracking performance, while the baseline drifts off the walking course and collides with a table.
To evaluate the HZD-CLF policy under real-world conditions, we deploy the robot outdoors across a variety of environments. As shown in Fig. 7, the test route includes diverse flat-ground surfaces such as concrete and tiled walkways, as well as mild slopes like ADA-compliant ramps. The robot traverses all terrain types within a single continuous 0.25-mile walking trial without any failures. The distance was determined by experimental design considerations, rather than reflecting any limitation of the policy.
@@
V Conclusion
We presented CLF-RL, a structured reward shaping framework that integrates reference trajectory tracking with control Lyapunov functions (CLFs). By embedding CLF-based objectives directly into the reinforcement learning reward, our approach replaces heuristic reward design with a theoretically grounded stability metric. We showed that incorporating the CLF decay condition improves tracking performance over baseline tracking-only rewards. CLF-based policies exhibit superior robustness across a range of perturbations in both simulation and hardware, demonstrating increased reliability.
Our framework is modular and extensible. We demonstrated its compatibility with both reduced-order (H-LIP) and full-order (HZD) reference generators, though other trajectory planners could be incorporated. While our training only used steady-state reference motions, the learned policies generalize well to transient motions—highlighting the framework’s flexibility. We validated the real-world robustness of the approach through extended outdoor trials, showing consistent and stable walking performance across varied flat-ground terrains. Finally, while our rewards are derived from CLFs that are theoretically sound for the full-order system, they are applied as guidance rather than hard constraints, leaving formal stability guarantees as a direction for future work.